OpenStudy (zyberg):

Two lubricating air heaters with 1 kW power and 220 V voltage are connected in series and turned on into 220 V voltage circuit. What is sum of the heaters' power then?

OpenStudy (zyberg):

I think that since in series circuit U = U_1 + U_2, and both heaters are the same we should have both of them using up 110 V? @ganeshie8

OpenStudy (zyberg):

Or maybe, since I = I_1 = I_2; P = UI; and P_1 = P_2; we find that P_1 / I _1 = P_2 / I_2; P / I = P / I; U_1 = U_2? and since U = U_1 + U_2; U = 2 * U_1; U_1 = U/2?

OpenStudy (zyberg):

(Basicially wrote the same thing as I had said in words, this time with letters, not sure if it's the right reasoning tho)

ganeshie8 (ganeshie8):

Your reasoning is good. Here is the trick : Find the resistance of the heaters

ganeshie8 (ganeshie8):

They gave you power and the voltage, can you find the resistance offered by the heater ?

OpenStudy (zyberg):

P = U^2/R; R = U^2 / P; R_1 = 220^2 / 1000 = 48.4 Ohms; R_2 = R_1.

ganeshie8 (ganeshie8):

Yes, next draw a rough diagram representing the circuit

OpenStudy (zyberg):